# Quantitative characterization of finite simple groups: a complement

### v3

@@ Acknowledgement
We sincerely thank the reviewers for their helpful suggestions for revisions. Ping Jin and Guohua Qian provided valuable advice on writing and specific content for this paper. Nanying Yang and Rulin Shen also offered their revision suggestions for this paper. In particular, Jinbao Li dedicated a considerable amount of effort to the editing and revising of this paper. The author would like to extend his heartfelt gratitude to all of them for their contributions.
Finally, the author would like to express his gratitude to Andreĭ Vasil’ev, Mariya Grechkoseeva and Natalia Maslova for their valuable suggestions on the revised version. The author thanks Xiaofang Gao for translating this article from Chinese version to revised version, which is helpful for the author’s communication at home and abroad.

### v5

@@ Acknowledgement
We sincerely thank the reviewers for their helpful suggestions for revisions. Ping Jin and Guohua Qian provided valuable advice on writing and specific content for this paper. Nanying Yang and Rulin Shen also offered their revision suggestions for this paper. In particular, Jinbao Li dedicated a considerable amount of effort to the editing and revising of this paper. The author would like to extend his heartfelt gratitude to all of them for their contributions. Finally, the author would like to express his gratitude to Andrey Vasil’ev, Mariya Grechkoseeva and Natalia Maslova for their valuable suggestions on the revised version. The author thanks Xiaofang Gao for translating this article from Chinese version to revised version, which is helpful for the author’s communication at home and abroad.
The author is very grateful to the referees for providing valuable revision suggestions and specific recommendations.